20220504_DX_Day. LBTI AO, Day 2

15:30 (UT) /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav - has a small bump at 565 (force +0.26)

15:44 (UT) bumping #565 - RB temperature 1.7C small bump at 565 (force +0.26)
  • + 50e-9 force +0.22
  • + 50e-9 force +0.18
  • + 50e-9 force +0.15
  • + 50e-9 force +0.11

15:44 UT RB temperature 2.1C bump at 565 (force +0.11, also 114 at +0.1)

15:44 UT saved flat: 20220504_154450/Ref_2C_flat.sav

15:50 set as default and reloaded

We found two issues:
  • amp_envelope.fits was missing in the KL_v28 directory
  • the FL fiber from LBTI was not plugged in the LBTI Switch BCU power #3 (was disconnected, instead the fiber dubbed ARGOS white/black pair, was plugged there)

18:52 UT applied flat: 20220504_154450/Ref_2C_flat.sav - RB temperature 3.5C
  • 565 (force +0.09), 114 (force +0.08)

18:55 UT first interaction matrix measurements are finally taking place

19:04 UT first AO CL with a low order reconstructor

-- GuidoBrusa - 04 May 2022
Topic revision: r3 - 05 May 2022, GuidoBrusa
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback